Luke Heckman has asked you to analyze the following investment alternatives to determine the highest after-tax rate of return under the assumption that he is subject to a 32 percent marginal federal income tax and a 5 percent state income tax (disregard the AMT for this problem). Which bond should he purchase?

•A corporate bond with a 7 percent pre-tax return.

•An out-of-state municipal bond with a 5.75 percent pre-tax return.

•An in-state municipal bond with a 5.5 percent pre-tax return.
Answer #1
Luke heckman
Corporate bond:
Interest on corporate bond is taxable at both levels i.e at state and federal.
state tax:- 7 * 5% = 0.35
post tax return = 7 - 0.35 = 6.65 %
Federal tax :- 6.65 * 32% = 2.128
Post tax return = 6.65 - 2.128 = 4.522%
Out of state municipal bond:
Interest on this type of bond is taxable at state level but not at federal level
state tax:- 5.75 * 5% = 0.2875
post tax return = 5.75 - 0.2875 = 5.4625 %
In state municipal bond:
Interest on this typeof bond is not taxable at any level.
post tax return = 5.5%
In state municipal bond should be purchased .